Please give a specific example. I understand what the addition rule is, however I am having trouble coming up with a situation in which it would be used while providing calculations. Also, please do not write in cursive as it is difficult for me to read. THANK YOU!

Describe a situation where you could use the addition rule to determine the probability of an outcome. Include an example of the calculations. (3 points)
Answer #1

Answer)

Addition rule states that, if the events A and B are mutually exclusive, that is nothing is common in between them,

P(A or B) = P(A) + P(B)

And if they are not mutually exclusive, then

P(A or B) = P(A) + P(B) - P(A and B)

Examples

For mutually exclusive:

You conducted a survey, about if people like dominos or kfc

Out of 700 people surveyed

300 said they like dominos

250 said they like kfc

And 150 said none

Probability of KFC or dominos

P(kfc) + p(dominos)

= (300/700) + (250/700)

= 0.7857142857142

If the events are not mutually exclusive:

You conducted survey, among women and men, about their favorite flavor of icream in between strawberry and vanilla

Total people surveyed = 700

400 women and 300 men

Out of 400 women, 250 said strawberry and 150 said vanilla.

And out of 300 men, 100 said strawberry and 200 said vanilla.

Now we need to find the probability that randomly selected person is women or likes vanilla

P(women) = 400/700

P(vanilla) = (150 + 200)/700

Now as there is common between women and vanilla which is 150, so we need to subtract it

P(women and vanilla) = 150/700

P(women or vanilla) =(400/700) + (350/700) - (150/700)

= 0.8571428571428